# A Bayesian approach for fitting semi-Markov mixture models of cancer latency to individual-level data

### v1

@@ 3.2 Partially observed data
Equation (7) indicates that the probability of positive screen $p^{i}_{j}$ is 00 when a person is in the $H$ compartment, while it becomes $\beta$ when they are in $P$; we see the parameter $\beta\in[0,1]$ is interpreted as the screening sensitivity and is not assumed to be known. Upon a cancer diagnosis—whether clinical or screen-detected—an individual receives treatment and leaves the screening program, implying that any positive screen is the last screen of the individual. Consequently, $p^{i}_{j}$ need not be defined when $t^{i}_{j}>\tau^{PC}_{i}$.
@@ 3.3 Bayesian inference via data-augmented Markov chain Monte Carlo
We now describe how to conduct Bayesian inference to fit the semi-Markov model defined by Equations (1)-(4) to censored data $\mathbf{Y}$. By Bayes Theorem, $\pi(\theta,\psi,\beta\mid\mathbf{Y})\propto\pi(\theta,\psi,\beta)L(\theta,\psi,\beta;\mathbf{Y})$ where $\pi(\theta,\psi,\beta)$ is the prior distribution of the parameters $\theta$, $\psi$ and $\beta$, and the marginal likelihood
[…]
is related to the complete-data likelihood (5) via integration. Equation (8) integrates over all the possible clinical histories of the individuals consistent with the observed data. No general closed-form expression for this integral exists in the semi-Markov case (Hsieh et al., 2002), and numerical integration over this high-dimensional latent space is intractable.
To overcome this difficult integration step, data-augmented MCMC takes advantage of the tractable form  (5). The observed data $\mathbf{Y}$ are augmented with the latent variables $\mathbf{Z}$ such that $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$ have closed-form expressions. We can then use Markov chain Monte Carlo to sample from the joint posterior distribution
@@
Given $M$ MCMC draws $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)\right\}_{m=1}^{M}$, marginalizing out latent variables a posteriori becomes trivial. The values $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)\right\}_{m=1}^{M}$ form an empirical approximation to the original distribution of interest $\pi\left(\theta,\psi,\beta\mid\mathbf{Y}\right)$. In other words, simply ignoring the latent variables in each posterior sample yields a valid marginal posterior of the parameters of interest. Similarly, $\left\{\mathbf{Z}^{(m)}\right\}_{m=1}^{M}$ form a Monte Carlo approximation to $\pi\left(\mathbf{Z}\mid\mathbf{Y}\right)$.
An advantage of the Bayesian formulation is that inference is not limited to $\theta$, $\psi$, $\beta$ and $\mathbf{Z}$. The posterior distribution $\pi\left(g(.)\mid\mathbf{Y}\right)$ of any functional $g(\theta,\psi,\beta,\mathbf{Z})$ is approximated by the transformed sample $\left\{g^{(m)}\right\}_{m=1}^{M}$, with $g^{(m)}=g\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)$. In Section 6, we will be interested in $g$ being the overdiagnosis rate.
In our model, the latent variables $\mathbf{Z}$ consist of the unknown transition ages into $P$ and the indolent status of each individual. We write $\mathbf{Z}=(\mathbf{Z}_{1},\dots,\mathbf{Z}_{n})$, where $\mathbf{Z}_{i}=\left(Z^{HP}_{i},Z^{I}_{i}\right)$ are the latent variables of individual $i$. $Z^{HP}_{i}\in(t_{0},c_{i}]\cup\varnothing$ denotes the pre-clinical onset age and $Z^{I}_{i}\in\{0,1\}$ the indolent status. Here, $\varnothing$ indicates the event that the corresponding transition age is larger than $c_{i}$.
Upon introducing these latent variables, we have closed-form expressions for $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$. First, by independence, we have
@@
$\tilde{L}_{i}(\theta,\psi;\mathbf{Z}_{i})\propto f(\mathbf{Z}_{i}\mid\theta,\psi)=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad Z^{HP}_{i}=\varnothing\\ f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)f_{\textsf{I}}(Z^{I}_{i};\psi),&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (11)
@@
$N_{i}(\theta,\psi,e_{i})$ $=\Pr(\tau^{PC}_{i}>e_{i}\mid\theta,\psi)$
$=\Pr(I_{i}=1)+\Pr\left(I_{i}=0,\tau^{HP}_{i}>e_{i}\right)+\Pr\left(I_{i}=0,\tau^{HP}_{i}<e_{i}<\tau^{PC}_{i}\right)$
$=\Pr(I_{i}=1)+\Pr(I_{i}=0)\left[\Pr\left(\tau^{HP}_{i}>e_{i}\mid I_{i}=0\right)+\Pr\left(\tau^{HP}_{i}<e_{i},e_{i}<\tau^{PC}_{i}\mid I_{i}=0\right)\right]$
[…]
scale the contribution from person $i$ by the probability that they did not develop clinical cancer before entering the study at age $e_{i}$. The inclusion of the factor (3.3) in the likelihood (10) comes from the fact that data from cohort analyses are left-truncated: study participants that developed clinical cancer before the study are excluded. Therefore, we condition on the fact that the individuals in the sample have not developed clinical cancer before entering the study. As long as $f_{\textsf{H}}$ and $S_{\textsf{prog}}$ are continuous and bounded, the univariate integral in (3.3) is easy to approximate numerically using a quadrature rule (e.g., the function integrate in base R).
@@
$f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})=\prod_{i=1}^{n}f(Y_{i}\mid\mathbf{Z}_{i},\theta,\psi,\beta),$ (13)
where
$f(Y_{i}\mid\mathbf{Z}_{i},\theta,\psi,\beta)=f_{t}\left(t^{PC}_{i}\mid\mathbf{Z}_{i},\theta\right)f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid\mathbf{Z}_{i},\beta\right);$ (14)
the distribution of the right-censored clinical onset age is
$f_{t}\left(t^{PC}_{i}\mid\mathbf{Z}_{i},\theta\right)=\begin{cases}\delta_{c_{i}}\left(t^{PC}_{i}\right),&\quad Z^{HP}_{i}=\varnothing\\ S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}\delta_{c_{i}}\left(t^{PC}_{i}\right)+f_{\textsf{prog}}\left(t^{PC}_{i}-Z^{HP}_{i};\theta\right)\delta_{\tau^{PC}_{i}}\left(t^{PC}_{i}\right),&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (15)
@@
where $m_{i}^{+}=\sum_{j=1}^{n_{i}}O^{i}_{j}$ denotes the number positive screens of person $i$, and $m_{i}^{-}=\sum_{j:Z^{HP}_{i}\leq t^{i}_{j}}\left(1-O^{i}_{j}\right)$ their number of false negative screens.
@@ 3.3.1 Prior specification
While the sampling methodology applies to any prior formulation $\pi(\theta,\psi,\beta)$, we assume for convenience here that the parameters are independent a priori
@@
Moreover, we specify the prior distribution of $\beta$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$ and that of $\psi$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$, where $\textsf{Be}(a,b)$ denotes to the beta distribution with density proportional to $x^{a-1}(1-x)^{b-1}$. The choice of a beta prior distribution for $\beta$ is motivated by conjugacy. Theorem 1 holds for any choice of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$.
Theorem 1 (Conjugacy).
@@ 4 Markov chain Monte Carlo scheme
We propose a data-augmented Markov chain Monte Carlo sampler that alternates between a Gibbs update for $\beta$, Metropolis-Hastings steps for the elements of $\theta$, and for $Z^{HP}_{1},\dots,Z^{HP}_{n}$, and a block Metropolis-Hastings update of $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$. Updating $\psi$ and $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)$ jointly is advantageous because we find these quantities are highly correlated a posteriori. This sampling scheme is summarized as a pseudo-algorithm in Algorithm 1. We note that methodologically, the elements of $\theta$ could also be jointly updated in a block Metropolis-Hastings step which may lead to further efficiency gains, though we do not explore this here.
Algorithm 1 Data-augmented Markov chain Monte Carlo
@@
     $\beta^{(m)}\sim\pi\left(\beta\mid\cdot\right)$  [Gibbs update, c.f. Theorem 1]
@@
         $\theta_{j}^{(m)}\leftarrow\text{Metropolis-Hasting}\left(\theta_{j};.\right)$  [update the $j$th element of $\theta$]
@@
         $\left(Z^{HP}_{i}\right)^{(m)}\leftarrow\text{Metropolis-Hasting}\left(Z^{HP}_{i};.\right)$
@@
     $\psi^{(m)},\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{(m)}\leftarrow\text{Metropolis-Hasting}\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n};.\right)$

### v2

@@ 33 $46$
Equation (5) indicates that the probability of positive screen $p^{i}_{j}$ is 0 when a person is in the $H$ compartment and is $\beta$ when they are in $P$; the parameter $\beta\in[0,1]$ is, therefore, interpreted as the screening sensitivity and is not assumed to be known. Upon a cancer diagnosis—whether clinical or screen-detected—an individual receives treatment and leaves the screening program, implying that any positive screen is the last screen of the individual. Consequently, $p^{i}_{j}$ need not be defined when $t^{i}_{j}>\tau^{PC}_{i}$.
@@ 3.3 Bayesian inference via data-augmented Markov chain Monte Carlo
We fit the semi-Markov model defined by Eqs. (1)-(5) to censored data $\mathbf{Y}$ in a Bayesian framework. By Bayes Theorem, $\pi(\theta,\psi,\beta\mid\mathbf{Y})\propto\pi(\theta,\psi,\beta)L(\theta,\psi,\beta;\mathbf{Y})$ where $\pi(\theta,\psi,\beta)$ is the prior on the parameters $(\theta,\psi,\beta)$, and the marginal likelihood
$L(\theta,\psi,\beta;\mathbf{Y})=\int f(\mathbf{Y}\mid\mathbf{x},\theta,\psi,\beta)L(\theta,\psi;\mathbf{x})d\mathbf{x},$ (6)
is related to the complete-data likelihood (4) via integration. Equation (6) integrates over all the possible clinical histories of the individuals consistent with the observed data. No general closed-form expression for this integral exists in the semi-Markov case 16, and numerical integration over this high-dimensional latent space is intractable.
To overcome this difficult integration step, DA-MCMC takes advantage of the tractable form  (4). The observed data $\mathbf{Y}$ are augmented with the latent variables $\mathbf{Z}$ such that $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$ have closed-form expressions. We can then use MCMC to sample from the joint posterior distribution
@@
Given $M$ MCMC draws $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)\right\}_{m=1}^{M}$, marginalizing out latent variables a posteriori becomes trivial. The values $\left\{\left(\theta^{(m)},\psi^{(m)},\beta^{(m)}\right)\right\}_{m=1}^{M}$ form an empirical approximation to the original distribution of interest $\pi\left(\theta,\psi,\beta\mid\mathbf{Y}\right)$. In other words, simply ignoring the latent variables in each posterior sample yields a valid marginal posterior of the parameters of interest. Similarly, $\left\{\mathbf{Z}^{(m)}\right\}_{m=1}^{M}$ form a Monte Carlo approximation to $\pi\left(\mathbf{Z}\mid\mathbf{Y}\right)$. An advantage of the Bayesian formulation is that inference is not limited to $\theta$, $\psi$, $\beta$ and $\mathbf{Z}$. The posterior distribution $\pi\left(g(.)\mid\mathbf{Y}\right)$ of any functional $g(\theta,\psi,\beta,\mathbf{Z})$ is approximated by the transformed sample $\left\{g^{(m)}\right\}_{m=1}^{M}$, with $g^{(m)}=g\left(\theta^{(m)},\psi^{(m)},\beta^{(m)},\mathbf{Z}^{(m)}\right)$. In Section 6, we will be let $g$ be the overdiagnosis rate.
In our model, the latent variables $\mathbf{Z}$ consist of the unobserved transition age into $P$ and indolent status of each individual. We write $\mathbf{Z}=(\mathbf{Z}_{1},\dots,\mathbf{Z}_{n})$, where $\mathbf{Z}_{i}=\left(Z^{HP}_{i},Z^{I}_{i}\right)$ are the latent variables of individual $i$. $Z^{HP}_{i}\in(t_{0},c_{i}]\cup c_{i}^{+}$ is the pre-clinical onset age and $Z^{I}_{i}\in\{0,1\}$ the indolent status. Here, $c_{i}^{+}$ indicates the event that the transition age is larger than $c_{i}$. Upon introducing these latent variables, we have closed-form expressions for $L(\theta,\psi;\mathbf{Z})$ and $f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})$. First, by independence, we have
@@
$\tilde{L}_{i}(\theta,\psi;\mathbf{Z}_{i})\propto f(\mathbf{Z}_{i}\mid\theta,\psi)=\begin{cases}S_{\textsf{H}}\left(c_{i}-t_{0};\theta\right),&\quad Z^{HP}_{i}=c_{i}^{+}\\ f_{\textsf{H}}\left(Z^{HP}_{i}-t_{0};\theta\right)f_{\textsf{I}}(Z^{I}_{i};\psi),&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (9)
@@
$N_{i}(\theta,\psi,e_{i}):=\Pr(\tau^{PC}_{i}>e_{i}\mid\theta,\psi)$ $=\Pr(I_{i}=1)+\Pr(I_{i}=0)\left[\Pr\left(\tau^{HP}_{i}>e_{i}\mid I_{i}=0\right)+\Pr\left(\tau^{HP}_{i}<e_{i},e_{i}<\tau^{PC}_{i}\mid I_{i}=0\right)\right]$
$=\psi+(1-\psi)\left[S_{\textsf{H}}\left(e_{i}-t_{0};\theta\right)+\int_{t_{0}}^{e_{i}}f_{\textsf{H}}(t-t_{0};\theta)S_{\textsf{prog}}\left(e_{i}-t;\theta\right)dt\right]$ (10)
scales the contribution from person $i$ by the probability that they did not develop clinical cancer before entering the study at age $e_{i}$. The inclusion of the factor (3.3) in the likelihood (8) comes from the fact that data from cohort analyses are left-truncated: participants that develop clinical cancer before the study started are excluded. Therefore, we condition on the fact that the individuals in the sample did not develop clinical cancer before the start of the study. As long as $f_{\textsf{H}}$ and $S_{\textsf{prog}}$ are continuous and bounded, the univariate integral in (3.3) is easy to approximate numerically using a quadrature rule (e.g., the function integrate in base R).
@@
$f(\mathbf{Y}\mid\theta,\psi,\beta,\mathbf{Z})=\prod_{i=1}^{n}f(Y_{i}\mid\mathbf{Z}_{i},\theta,\psi,\beta)=\prod_{i=1}^{n}f_{t}\left(t^{PC}_{i},\delta_{i}\mid\mathbf{Z}_{i},\theta\right)f_{O}\left(O^{i}_{1},\dots,O^{i}_{n_{i}}\mid\mathbf{Z}_{i},\beta\right),$ (11)
where the distribution of the right-censored clinical onset age is
$f_{t}\left(t^{PC}_{i},\delta_{i}\mid\mathbf{Z}_{i},\theta\right)=\begin{cases}1-\delta_{i},&\quad Z^{HP}_{i}=c_{i}^{+}\\ S_{\textsf{prog}}\left(c_{i}-Z^{HP}_{i};\theta\right)^{1-Z^{I}_{i}}(1-\delta_{i})+f_{\textsf{prog}}\left(t^{PC}_{i}-Z^{HP}_{i};\theta\right)\delta_{i},&\quad Z^{HP}_{i}\leq c_{i}\end{cases},$ (12)
@@
with $m_{i}^{+}=\sum_{j=1}^{n_{i}}O^{i}_{j}$ the number positive screens of person $i$, and $m_{i}^{-}=\sum_{j:Z^{HP}_{i}\leq t^{i}_{j}}\left(1-O^{i}_{j}\right)$ their number of false negative screens.
@@ 3.3.1 Prior specification
While the sampling methodology applies to any prior formulation $\pi(\theta,\psi,\beta)$, we assume for convenience that the parameters are independent a priori
@@
Moreover, we specify the prior distribution of $\beta$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$ and that of $\psi$ to be $\pi(\beta)=\textsf{Be}(a_{\beta},b_{\beta})$, where $\textsf{Be}(a,b)$ denotes to the beta distribution with density proportional to $x^{a-1}(1-x)^{b-1}$. The choice of a beta prior distribution for $\beta$ is motivated by conjugacy. Theorem 3.1 holds for any choice of $\textsf{F}_{\textsf{H}}$ and $\textsf{F}_{\textsf{prog}}$; its proof appears in web Appendix A.
Theorem 3.1 (Conjugacy).
@@ 4 Markov chain Monte Carlo scheme
We propose a DA-MCMC sampler that alternates between a Gibbs step for $\beta$, univariate Metropolis-Hastings steps for the elements of $\theta$, and for $Z^{HP}_{1},\dots,Z^{HP}_{n}$, and a block Metropolis-Hastings step of $\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n}\right)$ 27. This sampling scheme is summarized in Algorithm 1. The Gibbs step for $\beta$ proceeds by drawing a new value for this parameter from its full conditional distribution (16) given the most recent values of $\mathbf{Z}$. The proposal distribution for the Metropolis-Hastings steps are described in web Appendix B. Updating $\psi$ and $\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)$ jointly is crucial, because these quantities are highly correlated a posteriori. We note that methodologically, the elements of $\theta$ could also be jointly updated in a block Metropolis-Hastings step; though such block step may lead to further efficiency gains, we do not explore this here.
To initialize the DA-MCMC sampler, one only needs to specify the initial values of $(\theta,\psi,\beta)$. Given some initial values $\left(\theta^{(0)},\psi^{(0)},\beta^{(0)}\right)$, the algorithm generates the initial values of the latent variables $\mathbf{Z}^{(0)}$ using the independent proposal distributions of the Metropolis-Hastings steps described in web Appendix B.
@@
  $\beta^{(m)}\sim\pi\left(\beta\mid\cdot\right)$  [Gibbs update, c.f. Theorem 3.1]
@@
   $\theta_{j}^{(m)}\leftarrow\text{Metropolis-Hastings}\left(\theta_{j};.\right)$  [update the $j$th element of $\theta$]
@@
   $\left(Z^{HP}_{i}\right)^{(m)}\leftarrow\text{Metropolis-Hastings}\left(Z^{HP}_{i};.\right)$
@@
  $\psi^{(m)},\left(Z^{I}_{1},\dots,Z^{I}_{n}\right)^{(m)}\leftarrow\text{Metropolis-Hastings}\left(\psi,Z^{I}_{1},\dots,Z^{I}_{n};.\right)$